# Exploring the directly imaged HD 1160 system through spectroscopic characterisation and high-cadence variability monitoring

### v1

@@ Abstract
In this work, we further monitor and characterize HD 1160 B and its host star HD 1160 A using additional observations obtained with the Large Binocular Telescope (LBT). This includes a further night of variability monitoring of HD 1160 B with ALES+dgvAPP360, which we also use to produce its complementary spectral characterisation using the 2.8-4.2 m spectra from both epochs. We separately characterise the host star HD 1160 A using data obtained simultaneously with the PEPSI high resolution spectrograph. The differential spectrophotometry observations with ALES+dgvAPP360 also allow us to test the repeatability of the light curve precision achieved by Sutlieff et al. (2023) in their pilot study. In Section 2, we review the properties of the HD 1160 system. Our observations of this system are then described in Section 3, and in Section 4 we describe the methods used to reduce the data from each instrument and extract photometry of the targets. In Section 5, we investigate the variability of HD 1160 B by using the ALES+dgvAPP360 data to produce differential spectrophotometric light curves. We also use these data in Section 6 to produce and study the spectrum of HD 1160 B. The data obtained with PEPSI are analysed in Section 7, in which we explore the spectrum of HD 1160 A. The results found in each of these three sections are then discussed in Section 8. Finally, we summarise the conclusions of the paper in Section 9.
@@ 2 Target Properties
However, using observations from the Spectro-Polarimetric High-contrast imager for Exoplanets REsearch (SPHERE, Beuzit et al., 2019) instrument on the Very Large Telescope (VLT), Maire et al. (2016) concluded that the 1.0-1.6 m spectrum of HD 1160 B best matched that of a M6.0 dwarf. Unlike Nielsen et al. (2012), they did not find unusually red colours for either companion. They also found higher estimates for its mass; 79 MJup based on its luminosity and 107 MJup based on its effective temperature. The wide range of possible masses is driven by the uncertain age used, 100, which was chosen due to the lack of reliable age indicators with the upper limit given by the 300 Myr predicted main-sequence lifetime of an A0 star (Siess et al., 2000).
@@
The most recent spectral characterisation of HD 1160 B was carried out by Mesa et al. (2020), who again observed the system with SPHERE and found it to have a peculiar spectrum that is not well matched by any spectra in current spectral libraries, but concluded a spectral type of M5-M7 based on the best fits to individual spectral bands. They propose that this peculiarity could be explained by the presence of dust in its photosphere, or if it has a young age and is not yet fully matured. By fitting the spectrum of HD 1160 B with atmospheric models and considering alkali lines that become weaker at lower surface gravities, Mesa et al. (2020) found a low surface gravity of log(g) = 3.5-4.0 dex. This suggests that HD 1160 B may actually have a young age of 10-20 Myr, and a mass of 20 MJup, in contrast to previous results. However, they noted that they cannot rule out older ages.
@@ 4.1 LBT/ALES+dgvAPP360 data processing
Our goal is to use the LBT/ALES+dgvAPP360 observations to characterise HD 1160 B by measuring both its spectrum and its time variability. We therefore need to construct a flux-calibrated spectrum of the companion by summing the observations in the time dimension, and a ‘white-light’ curve of the companion by summing the observations in the wavelength dimension.
@@
As the data were obtained in pupil-stabilized mode, we could have applied post-processing algorithms reliant on angular diversity (e.g. Angular Differential Imaging, Marois et al., 2006) to further remove quasistatic speckle noise and increase the S/N of the targets. However, we chose not to do this so that we could make use of the stellar PSF provided by the dgvAPP360 as a simultaneous photometric reference when characterising the variability of HD 1160 B in Section 5. If we had applied an ADI-based algorithm the stellar PSF would have been removed, meaning there would be no photometric reference with which to divide out time-varying systematics from the companion flux.
@@ 4.2 PEPSI data processing
In this paper we aim to use the LBT/PEPSI observations to characterise the physical properties of host star HD 1160 A. The PEPSI data were reduced using the Spectroscopic Data Systems for PEPSI (SDS4PEPSI) generic package written in C++ under Linux and based on the 4A package for processing data from the SOFIN spectrograph on the Nordic Optical Telescope (Ilyin, 2000; Strassmeier et al., 2018a, b; Keles et al., 2022). SDS4PEPSI applied a fully automated set of standardised data reduction steps to the raw data, including CCD bias removal, photon noise estimation, flat-field correction, and scattered light subtraction. It then performed a weighted optimal extraction of the spectral orders to maximise the S/N of the target, and then performed wavelength calibration. The spectra were then normalised to the continuum by fitting the extracted spectral orders with a 2D smoothing spline on a regular grid of CCD pixels and echelle order numbers, and then shifted to the stellar rest frame. Each of these steps carried out by SDS4PEPSI is described in full detail by Strassmeier et al. (2018a). Finally, we combined the spectra from all of the exposures obtained with a given CD by interpolating them to the same wavelengths and summing them according to their weights, where the weights are defined as the inverse of the noise.
@@ 6 Spectral analysis of HD 1160 B
In addition to investigating the brightness fluctuations of HD 1160 B, we also extracted its spectrum on each night to allow us to characterise its physical properties through the fitting of atmospheric models.
@@ 6.2 Spectral fitting
Once we had obtained a flux calibrated spectrum of HD 1160 B for each night, we fit this data with atmospheric models to characterise its physical properties. We used a set of BT-Settl grid models (Allard et al., 2011, 2012, 2013) which were downloaded from the Spanish Virtual Observatory (SVO) Theory Server111http://svo2.cab.inta-csic.es/theory/newov2/. We restricted the models to those with effective temperatures between 400 K and 4600 K, surface gravities between 3.5 and 5.0 dex, metallicities between and 0.5, and an -enhancement of 0. The grid step sizes for temperature and surface gravity were 100 K and 0.5 dex, respectively, and metallicity could have values , 0, 0.3, or 0.5. We chose to restrict the range of possible surface gravities to these values based on the predicted physical limitations of objects with ages and masses within the ranges found for HD 1160 B in the recent literature, which are 10-125 Myr and 20 MJup to 123 MJup, respectively (Garcia et al., 2017; Curtis et al., 2019; Mesa et al., 2020). According to the isochrones and evolutionary tracks of the BT-Settl models, the surface gravities of objects with ages and masses within these constraints should always be 3.5 and 5.0 (Baraffe et al., 2015; Stone et al., 2016). Maire et al. (2016) did previously use a higher age upper limit of 300 Myr for the HD 1160 system, which would allow a HD 1160 B surface gravity of up to log(g) = 5.2, but Garcia et al. (2017) later found that such high ages were not consistent with the properties of the host star.
@@
Each model was convolved to the R40 spectral resolution of ALES and sampled at the wavelengths of our spectral data points. By fitting each model to the data, we then determined the scaling factor that minimises the Euclidean norm of the residual vector between the two i.e. the value multiplied by each model to best match it to the companion spectrum, and calculated the value for each fit accounting for the errors on each data point (e.g. Bohn et al., 2020a; Sutlieff et al., 2021). When calculating the values of model fits to high-contrast IFS data, it is important to consider the effects of spectral covariance arising from oversampling during the spectral extraction process and the wavelength-dependent behaviour of speckle noise (Greco & Brandt, 2016). We accounted for spectral covariances in the in the ALES IFS data by following the method described by Greco & Brandt (2016) to produce spectral covariance matrices for each night of data and apply them in our calculation. The model that produced the smallest value was then taken as the best-fitting model to the data. When performing this fitting procedure we excluded the data points in the shaded regions of Figure 9, which were not suitable for analysis as described in Section 6.1. We performed the fitting process three times; once each for the spectra from the first and second nights separately, and a third time considering both nights of data together. The best-fitting model in each case is shown overplotted on the companion spectrum in Figure 10. The best-fitting model to the first night of data alone has Teff = 2700 K, log(g) = 4.5 dex, and metallicity [Fe/H] = (purple line, Figure 10). When the second night of data is considered alone, the best-fitting model instead has solar metallicity and is slightly cooler, with Teff = 2300 K, log(g) = 5.0 dex (red line). This is likely due to the lower flux recorded in the 3-3.2 m region of the spectrum on this night. The effective temperature of the best-fitting model to both nights of data then lies between the two, as would be expected, with Teff = 2600 K, log(g) = 5.0 dex, and solar metallicity (green line). Using the values of each model fit as weights, we also calculated the weighted means and sided variance estimates (i.e. statistical errors) of these atmospheric parameters using the approach of Burgasser et al. (2010a, b); Stone et al. (2016). These results are presented in Table 3. The weighted means and their uncertainties are biased in some cases, where the preferred model fits lie at the edge of the allowed parameter range. We therefore instead report upper/lower limits in these instances.
@@
7 Characterising HD 1160 A with PEPSI
In addition to characterising HD 1160 B using the data obtained with ALES+dgvAPP360, the simultaneous high resolution PEPSI spectrum of the HD 1160 system in the optical (383-542 nm) further allows us to assess the properties of the host star HD 1160 A, which was originally classified as an A0 V star by Houk & Swift (1999). Although HD 1160 C lies at an angular separation far beyond the 2.25 diameter of the PEPSI fiber, HD 1160 B lies within this fiber diameter at a separation or 0.78, so the obtained PEPSI spectrum contains the spectra of both HD 1160 A and B. However, the contrast between the two is very large: mag in the 1.25 m J-band, and even larger at the shorter wavelengths covered by PEPSI (Garcia et al., 2017). We therefore assumed that the contribution of HD 1160 B to the PEPSI spectrum was negligible and treated the PEPSI spectrum as solely that of HD 1160 A (see Figure 11).
@@ 9 Conclusions
We present here a new study of the HD 1160 system using two nights of observations obtained with the Large Binocular Telescope. This work is divided into three parts: variability monitoring of red companion HD 1160 B with the dgvAPP360 coronagraph and the ALES IFS; a R40 spectral characterisation of HD 1160 B using the same data; and lastly a spectral characterisation of host star HD 1160 A using R = 50,000 high resolution spectroscopy obtained with the PEPSI spectrograph.
@@
We conducted our spectral characterisation of HD 1160 B by instead combining the LBT/ALES+dgvAPP360 observations over the time sequence for each night, thereby producing 2.9-4.1 m spectra of the companion. These spectra are the first for this target in the mid-infrared and are therefore highly complementary to previous studies in the literature. We find that the spectrum of HD 1160 B from the second night is systematically fainter in the 3.0-3.2 m wavelength range than on the first night, which could be due to the intrinsic variability of the companion if this difference is astrophysical. We then fit these spectra with BT-Settl atmospheric models, considering each night separately and both nights together, and found that the results differ considerably depending on the data being fitted. Our effective temperature Teff estimates range from 2279 K for the second night spectrum to 2794 K on the first night. This first night Teff is consistent with the literature, but those derived from the second night and both nights combined spectra are cooler. Our inferred luminosities are lower than those in the literature, but our radius estimates are mostly consistent. Overall, we conclude that the spectrum of HD 1160 B on the second night of our observations is not consistent with the literature. The differences in the results obtained for each spectrum highlights the impact that variability can have on atmospheric model fitting for substellar companions. Simultaneous observations over a broad wavelength range with facilities such as JWST may help to resolve the ambiguities arising from these model fits and determine whether the differences in the spectrum of HD 1160 B between epochs are due to time variability.
@@
Lastly, we performed a new characterisation of host star HD 1160 A by comparing the R50,000 high resolution spectrum obtained with PEPSI to BT-NextGen atmospheric models. We found values for the physical properties of HD 1160 A; Teff = 9200 K, log(g) = 3.5, and = 96 km s-1. This model corresponds to a spectral type of A1 IV-V, which is slightly later than the literature A0 V classification found by Houk & Swift (1999) using photographic plates. This may explain the apparent underluminosity of HD 1160 A previously noted by Nielsen et al. (2012). By considering our rotational velocity measurement alongside the known near edge-on inclination angle of the HD 1160 system, we find that HD 1160 A rotates slower than the typical A-type star, and hence rule out gravitational darkening as the cause of any underluminosity.

### v2

# Exploring the directly imaged HD 1160 system through spectroscopic characterization and high-cadence variability monitoring

@@ Abstract
In this work, we further monitor and characterize HD 1160 B and its host star HD 1160 A using additional observations obtained with the Large Binocular Telescope (LBT). This includes a further night of variability monitoring of HD 1160 B with ALES+dgvAPP360, which we also use to produce its complementary spectral characterization using the 2.8-4.2 m spectra from both epochs. We separately characterize the host star HD 1160 A using data obtained simultaneously with the PEPSI high resolution spectrograph. The differential spectrophotometry observations with ALES+dgvAPP360 also allow us to test the repeatability of the light curve precision achieved by Sutlieff et al. (2023) in their pilot study. In Section 2, we review the properties of the HD 1160 system. Our observations of this system are then described in Section 3, and in Section 4 we describe the methods used to reduce the data from each instrument and extract photometry of the targets. In Section 5, we investigate the variability of HD 1160 B by using the ALES+dgvAPP360 data to produce differential spectrophotometric light curves. We also use these data in Section 6 to produce and study the spectrum of HD 1160 B. The data obtained with PEPSI are analysed in Section 7, in which we explore the spectrum of HD 1160 A. The results found in each of these three sections are then discussed in Section 8. Finally, we summarise the conclusions of the paper in Section 9.
@@ 2 Target Properties
However, using observations from the Spectro-Polarimetric High-contrast imager for Exoplanets REsearch (SPHERE, Beuzit et al., 2019) instrument on the Very Large Telescope (VLT), Maire et al. (2016) concluded that the 1.0-1.6 m spectrum of HD 1160 B best matched that of a M6.0 dwarf. Unlike Nielsen et al. (2012), they did not find unusually red colours for either companion. They also found higher estimates for its mass; 79 MJup based on its luminosity and 107 MJup based on its effective temperature. The wide range of possible masses is driven by the uncertain age used, 100 Myr, which was chosen due to the lack of reliable age indicators with the upper limit given by the 300 Myr predicted main-sequence lifetime of an A0 star (Siess et al., 2000).
@@
The most recent spectral characterization of HD 1160 B was carried out by Mesa et al. (2020), who again observed the system with SPHERE and found it to have a peculiar spectrum that is not well matched by any spectra in current spectral libraries, but concluded a spectral type of M5-M7 based on the best fits to individual spectral bands. They propose that this peculiarity could be explained by the presence of dust in its photosphere, or if it has a young age and is not yet fully matured. By fitting the spectrum of HD 1160 B with atmospheric models and considering alkali lines that become weaker at lower surface gravities, Mesa et al. (2020) found a low surface gravity of log(g) = 3.5-4.0 dex. This suggests that HD 1160 B may actually have a young age of 10-20 Myr, and a mass of 20 MJup, in contrast to previous results. However, they noted that they cannot rule out older ages.
@@ 4.1 LBT/ALES+dgvAPP360 data processing
Our goal is to use the LBT/ALES+dgvAPP360 observations to characterize HD 1160 B by measuring both its spectrum and its time variability. We therefore need to construct a flux-calibrated spectrum of the companion by summing the observations in the time dimension, and a ‘white-light’ curve of the companion by summing the observations in the wavelength dimension.
@@
As the data were obtained in pupil-stabilized mode, we could have applied post-processing algorithms reliant on angular diversity (e.g. Angular Differential Imaging, Marois et al., 2006) to further remove quasistatic speckle noise and increase the S/N of the targets. However, we chose not to do this so that we could make use of the stellar PSF provided by the dgvAPP360 as a simultaneous photometric reference when characterizing the variability of HD 1160 B in Section 5. If we had applied an ADI-based algorithm the stellar PSF would have been removed, meaning there would be no photometric reference with which to divide out time-varying systematics from the companion flux.
@@ 4.2 PEPSI data processing
In this paper we aim to use the LBT/PEPSI observations to characterize the physical properties of host star HD 1160 A. The PEPSI data were reduced using the Spectroscopic Data Systems for PEPSI (SDS4PEPSI) generic package written in C++ under Linux and based on the 4A package for processing data from the SOFIN spectrograph on the Nordic Optical Telescope (Ilyin, 2000; Strassmeier et al., 2018a, b; Keles et al., 2022). SDS4PEPSI applied a fully automated set of standardised data reduction steps to the raw data, including CCD bias removal, photon noise estimation, flat-field correction, and scattered light subtraction. It then performed a weighted optimal extraction of the spectral orders to maximise the S/N of the target, and then performed wavelength calibration. The spectra were then normalised to the continuum by fitting the extracted spectral orders with a 2D smoothing spline on a regular grid of CCD pixels and echelle order numbers, and then shifted to the stellar rest frame. Each of these steps carried out by SDS4PEPSI is described in full detail by Strassmeier et al. (2018a). Finally, we combined the spectra from all of the exposures obtained with a given CD by interpolating them to the same wavelengths and summing them according to their weights, where the weights are defined as the inverse of the noise.
@@ 6 Spectral analysis of HD 1160 B
In addition to investigating the brightness fluctuations of HD 1160 B, we also extracted its spectrum on each night to allow us to characterize its physical properties through the fitting of atmospheric models.
@@ 6.2 Spectral fitting
Once we had obtained a flux calibrated spectrum of HD 1160 B for each night, we fit this data with atmospheric models to characterize its physical properties. We used a set of BT-Settl grid models (Allard et al., 2011, 2012, 2013) which were downloaded from the Spanish Virtual Observatory (SVO) Theory Server111http://svo2.cab.inta-csic.es/theory/newov2/. We restricted the models to those with effective temperatures between 400 K and 4600 K, surface gravities between 3.5 and 5.0 dex, metallicities between and 0.5, and an -enhancement of 0. The grid step sizes for temperature and surface gravity were 100 K and 0.5 dex, respectively, and metallicity could have values , 0, 0.3, or 0.5. We chose to restrict the range of possible surface gravities to these values based on the predicted physical limitations of objects with ages and masses within the ranges found for HD 1160 B in the recent literature, which are 10-125 Myr and 20 MJup to 123 MJup, respectively (Garcia et al., 2017; Curtis et al., 2019; Mesa et al., 2020). According to the isochrones and evolutionary tracks of the BT-Settl models, the surface gravities of objects with ages and masses within these constraints should always be 3.5 and 5.0 (Baraffe et al., 2015; Stone et al., 2016). Maire et al. (2016) did previously use a higher age upper limit of 300 Myr for the HD 1160 system, which would allow a HD 1160 B surface gravity of up to log(g) = 5.2, but Garcia et al. (2017) later found that such high ages were not consistent with the properties of the host star.
@@
Each model was convolved to the R40 spectral resolution of ALES and sampled at the wavelengths of our spectral data points. By fitting each model to the data, we then determined the scaling factor that minimises the Euclidean norm of the residual vector between the two i.e. the value multiplied by each model to best match it to the companion spectrum, and calculated the value for each fit accounting for the errors on each data point (e.g. Bohn et al., 2020a; Sutlieff et al., 2021). When calculating the values of model fits to high-contrast IFS data, it is important to consider the effects of spectral covariance arising from oversampling during the spectral extraction process and the wavelength-dependent behaviour of speckle noise (Greco & Brandt, 2016). We accounted for spectral covariances in the ALES IFS data by following the method described by Greco & Brandt (2016) to produce spectral covariance matrices for each night of data and apply them in our calculation. The model that produced the smallest value was then taken as the best-fitting model to the data. When performing this fitting procedure we excluded the data points in the shaded regions of Figure 9, which were not suitable for analysis as described in Section 6.1. We performed the fitting process three times; once each for the spectra from the first and second nights separately, and a third time considering both nights of data together. The best-fitting model in each case is shown overplotted on the companion spectrum in Figure 10. The best-fitting model to the first night of data alone has Teff = 2700 K, log(g) = 4.5 dex, and metallicity [Fe/H] = (purple line, Figure 10). When the second night of data is considered alone, the best-fitting model instead has solar metallicity and is slightly cooler, with Teff = 2300 K, log(g) = 5.0 dex (red line). This is likely due to the lower flux recorded in the 3-3.2 m region of the spectrum on this night. The effective temperature of the best-fitting model to both nights of data then lies between the two, as would be expected, with Teff = 2600 K, log(g) = 5.0 dex, and solar metallicity (green line). Using the values of each model fit as weights, we also calculated the weighted means and sided variance estimates (i.e. statistical errors) of these atmospheric parameters using the approach of Burgasser et al. (2010a, b); Stone et al. (2016). These results are presented in Table 3. The weighted means and their uncertainties are biased in some cases, where the preferred model fits lie at the edge of the allowed parameter range. We therefore instead report upper/lower limits in these instances.
@@
7 Characterizing HD 1160 A with PEPSI
In addition to characterizing HD 1160 B using the data obtained with ALES+dgvAPP360, the simultaneous high resolution PEPSI spectrum of the HD 1160 system in the optical (383-542 nm) further allows us to assess the properties of the host star HD 1160 A, which was originally classified as an A0 V star by Houk & Swift (1999). Although HD 1160 C lies at an angular separation far beyond the 2.25 diameter of the PEPSI fiber, HD 1160 B lies within this fiber diameter at a separation or 0.78, so the obtained PEPSI spectrum contains the spectra of both HD 1160 A and B. However, the contrast between the two is very large: mag in the 1.25 m J-band, and even larger at the shorter wavelengths covered by PEPSI (Garcia et al., 2017). We therefore assumed that the contribution of HD 1160 B to the PEPSI spectrum was negligible and treated the PEPSI spectrum as solely that of HD 1160 A (see Figure 11).
@@ 9 Conclusions
We present here a new study of the HD 1160 system using two nights of observations obtained with the Large Binocular Telescope. This work is divided into three parts: variability monitoring of red companion HD 1160 B with the dgvAPP360 coronagraph and the ALES IFS; a R40 spectral characterization of HD 1160 B using the same data; and lastly a spectral characterization of host star HD 1160 A using R = 50,000 high resolution spectroscopy obtained with the PEPSI spectrograph.
@@
We conducted our spectral characterization of HD 1160 B by instead combining the LBT/ALES+dgvAPP360 observations over the time sequence for each night, thereby producing 2.9-4.1 m spectra of the companion. These spectra are the first for this target in the mid-infrared and are therefore highly complementary to previous studies in the literature. We find that the spectrum of HD 1160 B from the second night is systematically fainter in the 3.0-3.2 m wavelength range than on the first night, which could be due to the intrinsic variability of the companion if this difference is astrophysical. We then fit these spectra with BT-Settl atmospheric models, considering each night separately and both nights together, and found that the results differ considerably depending on the data being fitted. Our effective temperature Teff estimates range from 2279 K for the second night spectrum to 2794 K on the first night. This first night Teff is consistent with the literature, but those derived from the second night and both nights combined spectra are cooler. Our inferred luminosities are lower than those in the literature, but our radius estimates are mostly consistent. Overall, we conclude that the spectrum of HD 1160 B on the second night of our observations is not consistent with the literature. The differences in the results obtained for each spectrum highlights the impact that variability can have on atmospheric model fitting for substellar companions. Simultaneous observations over a broad wavelength range with facilities such as JWST may help to resolve the ambiguities arising from these model fits and determine whether the differences in the spectrum of HD 1160 B between epochs are due to time variability.
@@
Lastly, we performed a new characterization of host star HD 1160 A by comparing the R50,000 high resolution spectrum obtained with PEPSI to BT-NextGen atmospheric models. We found values for the physical properties of HD 1160 A; Teff = 9200 K, log(g) = 3.5, and = 96 km s-1. This model corresponds to a spectral type of A1 IV-V, which is slightly later than the literature A0 V classification found by Houk & Swift (1999) using photographic plates. This may explain the apparent underluminosity of HD 1160 A previously noted by Nielsen et al. (2012). By considering our rotational velocity measurement alongside the known near edge-on inclination angle of the HD 1160 system, we find that HD 1160 A rotates slower than the typical A-type star, and hence rule out gravitational darkening as the cause of any underluminosity.